a train leaves a station and travels east at 72km/h. three hours later a second train leaves on a parallel track and travels at 120km/h. When will the second train overtake the first train?
Answer by ptaylor(2198)


Distance(d)=Rate(r) times Time(t) or d=rt; t=d/r and r=d/t
Let t=number of hours that elapses after the second train leaves until the first train is overtaken
When the second train leaves, the first train has already
travelled 72*3 or 216 mi
Now we know that when the two trains have travelled the same distance, the second train will have overtaken the first.
distance first train travels=216+72t
distance 2nd train travels=120t
So. our equation to solve is:
216+72t=120t subtract 72t from both sides
216+72t-72t=120t-72t collect like terms
216=48t divide both sides by 48
t=4.5 hours-----------------------time it takes 2nd train to overtake the first train after the 2nd train leaves.
CK
216+72*4.5=120*4.5
540=540

Hope this helps---ptaylor